OpenStudy (anonymous):

what are the zeros of P(x)=-3(x-2)(x+1)(x+i)(x-i)

OpenStudy (anonymous):

Umm.... did you do a mistake with i?

OpenStudy (anonymous):

no, those are imaginary

hero (hero):

Hint: set each factor equal to zero, then solve for x: x - 2 = 0 x + 1 = 0 x + i = 0 x - i = 0

OpenStudy (anonymous):

thanks, i got that much.. but wasn't so sure